# CFBench: A Comprehensive Constraints-Following Benchmark for LLMs

### v1

@@ Abstract
The adeptness of Large Language Models (LLMs) in comprehending and following natural language instructions is critical for their deployment in sophisticated real-world applications. Existing evaluations mainly focus on fragmented constraints or narrow scenarios, but they overlook the comprehensiveness and authenticity of constraints from the user’s perspective. To bridge this gap, we propose CFBench, a large-scale Comprehensive Constraints Following Benchmark for LLMs, featuring 1,000 curated samples that cover more than 200 real-life scenarios and over 50 NLP tasks. CFBench meticulously compiles constraints from real-world instructions and constructs an innovative systematic framework for constraint types, which includes 10 primary categories and over 25 subcategories, and ensures each constraint is seamlessly integrated within the instructions. To make certain that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization, covering various perspectives of constraints, instructions, and requirement fulfillment. Evaluating current leading LLMs on CFBench reveals substantial room for improvement in constraints following, and we further investigate influencing factors and enhancement strategies. The data and code are publicly available at https://github.com/PKU-Baichuan-MLSystemLab/CFBench.
@@ Introduction
Large Language Models (LLMs) have become the cornerstone of numerous cutting-edge research tasks and are widely utilized in real-world scenarios (Brown et al. 2020; Chowdhery et al. 2023; Achiam et al. 2023; Touvron et al. 2023). In real-world scenarios, human instructions are inherently complex and accompanied by explicit constraints, requiring models to both understand intricate requirements and strictly comply with these constraints (Yang et al. 2023; Zhong et al. 2021; Mishra et al. 2022; Wei et al. 2021; Sanh et al. 2022). The proficiency of LLMs in comprehending requirements and adhering to natural language constraints is essential, as it ensures tasks are executed precisely and resolved perfectly according to user instructions.
@@
The prevailing method for evaluating a model’s instruction-following ability involves using quantitative programs, human evaluators, or advanced LLMs to assess response quality across single constraints, complex problems, and finite constraints  (Zhou et al. 2023a; Wang et al. 2023; Li et al. 2023; Zheng et al. 2024; Xu et al. 2023).  Laskar et al. (2024) underscores the importance of evaluating data quality, highlighting the necessity for real and extensive data distribution, along with its applicability to real-world scenarios.  Sun et al. (2024b) also stresses that realistic evaluation metrics reflect model capabilities and guide iteration. Constraints-following evaluation faces analogous challenges, particularly within complex real-world scenarios, where data sources and contexts are diverse, and where evaluation is both subjective and arduous. Fig. 1, which addresses the aforementioned challenges, presents a sample from CFBench illustrating the Trump assassination event with different colors representing various constraint types. The instruction include multiple constraints, and the evaluation method uses a checklist to break down complex requirements into independent checkpoints, annotating constraint types and priorities. LLMs are then used to assess each checkpoint. Consequently, we introduce two more profound challenges in constraints-following assessment.
Q1: How to construct high-quality evaluation data?
@@
Many studies focus on evaluating single constraint (Chen et al. 2022; Tang et al. 2023), lacking comprehensive analysis across diverse constraints.  He et al. (2024b) examines LLM performance on complex real-world instructions but neglect constraint diversity and scenario coverage. Jiang et al. (2023) incrementally incorporate fine-grained constraints to craft multi-level instructions. However, with only 75 instances of mixed type, which risks variability due to limited data, and equating difficulty with constraint quantity oversimplifies the task. Recent work focuses on evaluating constraints combinability (Wen et al. 2024). To ensure data quality, we systematically categorize constraints by mining real-world online data and using classification, synthesis, and expert design, covering 10 primary categories and over 25 subcategories. We also cross-match these constraints with various domains and scenarios, ensuring balanced representation and expert-validated reasonableness.
Q2:How to evaluate accurately and meticulously?
@@
Evaluating LLMs’ adherence to constraints is challenging and typically involves manual, automated, and programmatic assessments using various metrics. Representative work computes outcomes for verifiable instructions using code  (Zhou et al. 2023a; He et al. 2024b).  Jiang et al. (2023) uses scripts and constraint-evolution paths to handle diverse challenging instructions, introducing three metrics tailored to the data’s characteristics. The DRFR method decomposes complex constraints into binary judgments, with GPT evaluating each criterion (Qin et al. 2024). Indeed, previous work has ensured the feasibility and objectivity of evaluations through various methods, but they have overlooked assessments from the user’s multiple perspectives. We deconstruct complex instructions from the user’s perspective into multiple sub-needs, categorizing them by priority and constraint type, with LLMs evaluating each checkpoint. Furthermore, a multi-dimensional evaluation criteria is proposed using three metrics from the perspectives of constraints, instructions, and requirements priority.
@@
We introduce CFBench, a comprehensive benchmark designed to thoroughly evaluate the constraint comprehension and following capabilities of LLMs. CFBench aggregates constraints from real-world scenarios and pioneers a hierarchical system for constraint types, encompassing 10 primary categories and over 25 secondary subcategories organized through taxonomic and statistical methodologies. CFBench features 1,000 meticulously curated samples spanning more than 200 real-life scenarios across 20 domains and over 50 NLP tasks, enhancing the breadth and generality of the evaluation data. Additionally, we have seamlessly integrated original instructions and constraint types within each sample, paying particular attention to nuanced combinations, ensuring each constraint is credibly and coherently embedded. Our advanced evaluation methodology incorporates multi-dimensional assessment criteria, which prioritizing requirements to align LLM outputs with user needs, enhance interpretability, and facilitate iterative development. Finally, extensive experiments and exploratory discussions provide strong support for evaluation and optimization.
@@
  • We introduce CFBench, a meticulously annotated, large-scale, high-quality benchmark that encompasses a broad spectrum of real-world scenarios and NLP tasks.
@@
Figure 2: The construction pipeline of CFBench. Initially, it entails the construction of the constraint system, followed by the assembly of the dataset, and culminating in the proposal of a multi-perspective user view evaluation.
Related Work
Evaluation for LLMs
Numerous studies have concurrently evaluated LLMs from various perspectives. By integrating a multitude of existing datasets, many have assessed the overall capabilities of LLMs  (Bommasani, Liang, and Lee 2023; Zhong et al. 2024; Dubois et al. 2024; Chia et al. 2024; Hendrycks et al. 2021). Some research has delved into specialized capabilities, such as programming  (Chen et al. 2021), reasoning  (Cobbe et al. 2021), and knowledge  (Huang et al. 2024). Unlike previous studies, we evaluate LLMs’ instruction-following abilities with a focus on constraint-based instructions.
Instruction Following
Fine-tuning LLMs with annotated instructional data enhances their ability to follow general language instructions  (Weller et al. 2020; Sanh et al. 2022; Mishra et al. 2022). Studies show that more complex or constrained instructions further improve this ability. For instance, six methods to create intricate instructions from a small set of handwritten seed data are proposed (Xu et al. 2023; Luo et al. 2023), while  Mukherjee et al. (2023) elevate training data complexity by having GPT-4 generate reasoning steps for simple instructions. The latest work  (Sun et al. 2024a; He et al. 2024a; Dong et al. 2024) suggests that increasing the number and variety of constraints can enhance the complexity of instructions, thereby further improving the model’s ability to follow constraint-based instructions.
Evaluation of Constraints Following
Constraints such as word count, position, topics, and content have garnered significant attention in the field of Controlled Text Generation  (Yao et al. 2023; Zhou et al. 2023b; Chen et al. 2022). Zhou et al. (2023a) centers on assessing 25 verifiable instructions. Numerous studies have explored the adherence of LLMs to format constraints, such as complex tabular data  (Tang et al. 2023) and customized scenario formats  (Xia et al. 2024).  Qin et al. (2024) decomposing a single instruction into multiple constraints. He et al. (2024b) gathered constraints from real-world scenarios and developed a sophisticated benchmark using detailed task descriptions and inputs.  Jiang et al. (2023) progressively integrates fine-grained constraints to develop multi-level instructions, thereby enhancing complexity across six distinct types. Concurrent work  (Wen et al. 2024), constructs a novel benchmark by synthesizing and refining data from the aforementioned benchmarks, with an emphasis on the combinatorial types of constraints. However, previous studies suffered from fragmented constraints, limited scenarios, and misaligned evaluation methods with user perspectives.
@@
Constraints System
Constraints Collection
We amass a diverse corpus of instructions from real-world scenarios and various NLP tasks  (Xia et al. 2024; Li et al. 2024) to ensure a comprehensive system. Initially, we aggregate several million instructions from online logs and NLP tasks, refining these through length filtering and clustering to distill 30,000 high-quality instructions. Utilizing advanced LLM techniques, we extract and expand atomic constraints through evolutionary methods. Using LLMs, we carefully select meaningful atomic constraints, resulting in over 5000 unique constraints. Domain experts first filter out unreasonable or meaningless atomic constraints and then synthesize these into a structured framework with 10 primary categories and 25 subcategories, guided by principles of statistics, taxonomy, and linguistics.
Constraints System
Content constraints
control the scope and depth of output content by specifying certain conditions (Zhang et al. 2023), and can be divided into lexical constraints, element constraints, and semantic constraints based on their granularity. Numerical constraints, which ensure that output content meets length and quantity requirements (Yao et al. 2023), can be classified into word-level, sentence-level, paragraph-level, document-level based on the objects involved in the planning. Stylistic constraints impart a unique flavor and color to the output, revealing the author’s traits and chosen social objectives (Tsai et al. 2021), can be subdivided into tonal, formal, audience, and authorial style constraints based on the perspective of application. Format constraints (Tang et al. 2023) standardize expression to guide the generation of complex content and can be categorized into fundamental, bespoke, and specialized scenario constraints based on their usage scenarios. Linguistic constraints (Zhou et al. 2023b) adapt to various scenarios by controlling internal features and logic, grouped into Pragmatic, Syntactic, Morphological, Phonological, and other constraints. Situation constraints (Liu et al. 2023) guide response appropriateness through background or situational parameters, can be classified into role-based, task-specific, and complex contextual constraints. Example constraints regulate new responses by leveraging the intrinsic patterns established by a limited set of samples, with an emphasis on assessing the model’s proficiency in contextual constraint learning. Inverse constraints narrow the response space through the mechanism of indirect exclusion. Contradictory constraints denote conditions that are mutually exclusive, rendering it impossible for the response to fulfill all requirements concurrently, which are prevalent in online logs and are often easily overlooked. Rule constraints define logic flows or actions and meticulously crafted to standardize the road of responses. Details are in Appendix Tab. 8.
Dataset Construction
@@
Data Source and Selection
Real-world scenarios and NLP tasks form the foundation for CFBench’s initial instructions. By harnessing advanced LLMs, we rigorously assess each instruction for constraint types and quantities within a predefined system, filtering out those with unreasonable or ineffective constraints. Subsequently, we balance the scenarios and constraint types, resulting in a refined set of 2,000 instructions covering all scenarios and NLP tasks. Prompts for instruction and checklist generation are in Appendix.
Iterative Refinement
Professional annotators carefully review and refine the data, ensuring the rationality of constraints and gold answer. If modifications are needed, instructions are revised, and LLMs generate responses with refined evaluation criteria, repeating this process until satisfactory results are achieved. Ultimately, comprehensive support is formulated for each sample, detailing high-quality instructions, the ideal answer, specific assessment criteria, constraint types, and priority levels, totaling 1000 entries.
@@
Dataset Statistics
Overall Statistics
Tab. 1 provides a statistical overview of CFBench, highlighting the substantial differences between the two sets. The Hard Set features more detailed instructions, a greater variety of constraints, and a higher number of constraints, indicating a higher level of complexity compared to the Easy Set. Furthermore, the table demonstrates the diversity and balanced distribution of primary constraint types within CFBench, an area where we significantly outperform other benchmarks. See the Appendix for division details.
Task and Domain Distribution
@@
Comparison with Other Benchmarks
As shown in Tab. 3, we thoroughly compare our benchmark with various relevant ones. In terms of size, our benchmark contains approximately twice the number of samples as others. Notably, FollowBench (Jiang et al. 2023), despite having 820 samples, originally had only one-fifth of this number, with each sample expanded by adding five constraints. From the constraint perspective, CFBench is the only benchmark with systematized constraints, featuring the highest number of primary constraint types and ensuring a balanced consideration of these types. Regarding evaluation, InFoBench (Qin et al. 2024) also sets response criteria for each sample, while CFBench distinguishes itself by incorporating user demand prioritization and emphasizing objectivity.
Evaluation Protocol
Evaluation Criteria
We breaking down instructions into multiple simple, independent checkpoints to ensure evaluation accuracy, inspiration was drawn from DRFR (Qin et al. 2024). This approach addresses the challenge of evaluating entire responses, especially for complex instructions with multiple constraints. Unlike DRFR, our method emphasizes defining ideal response characteristics and critical evaluation points. The previous sections detailed the checklist generation process, a key part of our evaluation criteria. Furthermore, we employ the state-of-the-art LLM, GPT-4o, as the evaluation model. By repeatedly feeding it the instruction, test model response, and checklist with a carefully tuned prompt, we ensure that the judged response fully meets the judgement format check. This iterative process aims to maximize confidence in our evaluation. The specific evaluation prompt is in the Appendix.
Evaluation Metrics

### v2

@@ Abstract
The adeptness of Large Language Models (LLMs) in comprehending and following natural language instructions is critical for their deployment in sophisticated real-world applications. Existing evaluations mainly focus on fragmented constraints or narrow scenarios, but they overlook the comprehensiveness and authenticity of constraints from the user’s perspective. To bridge this gap, we propose CFBench, a large-scale Chinese Comprehensive Constraints Following Benchmark for LLMs, featuring 1,000 curated samples that cover more than 200 real-life scenarios and over 50 NLP tasks. CFBench meticulously compiles constraints from real-world instructions and constructs an innovative systematic framework for constraint types, which includes 10 primary categories and over 25 subcategories, and ensures each constraint is seamlessly integrated within the instructions. To make certain that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization, covering various perspectives of constraints, instructions, and requirement fulfillment. Evaluating current leading LLMs on CFBench reveals substantial room for improvement in constraints following, and we further investigate influencing factors and enhancement strategies. The data and code will be made available.
CFBench: A Comprehensive Constraints-Following Benchmark for LLMs
Tao Zhang1thanks:  Equal contribution.  Chenglin Zhu211footnotemark: 1  Yanjun Shen [^1] footnotemark: 1  Wenjing Luo1  Yan Zhang1 Hao Liang2Tao Zhang1Fan Yang1Mingan Lin1Yujing Qiao1 Weipeng Chen1Bin Cui2Wentao Zhang2thanks:  Corresponding author.Zenan Zhou122footnotemark: 2 1Baichuan Inc.  2Peking University {zhangtao2, shenyanjun, zhouzenan}@baichuan-inc.com zhuchenglin@stu.pku.edu.cn, wentao.zhang@pku.edu.cn
1 Introduction
@@
Large Language Models (LLMs) have become the cornerstone of numerous cutting-edge research tasks and are widely utilized in real-world scenarios (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2023; Touvron et al., 2023). In real-world scenarios, human instructions are inherently complex and accompanied by explicit constraints, requiring models to both understand intricate requirements and strictly comply with these constraints Yang et al. (2023); Zhong et al. (2021); Mishra et al. (2022); Wei et al. (2021); Sanh et al. (2022). The proficiency of LLMs in comprehending requirements and adhering to natural language constraints is essential, as it ensures tasks are executed precisely and resolved perfectly according to user instructions.
@@
The prevailing method for evaluating a model’s instruction-following ability involves using quantitative programs, human evaluators, or advanced LLMs to assess response quality across single constraints, complex problems, and finite constraints  Zhou et al. (2023a); Wang et al. (2023); Li et al. (2023); Zheng et al. (2024); Xu et al. (2023).  Laskar et al. (2024) underscores the importance of evaluating data quality, highlighting the necessity for real and extensive data distribution, along with its applicability to real-world scenarios.  Sun et al. (2024b) also stresses that realistic evaluation metrics reflect model capabilities and guide iteration. Constraints-following evaluation faces analogous challenges, particularly within complex real-world scenarios, where data sources and contexts are diverse, and where evaluation is both subjective and arduous. Fig. 1, which addresses the aforementioned challenges, presents a sample from CFBench illustrating the Trump assassination event with different colors representing various constraint types. The instruction include multiple constraints, and the evaluation method uses a checklist to break down complex requirements into independent checkpoints, annotating constraint types and priorities. LLMs are then used to assess each checkpoint. For the English-Chinese comparison example, see Appendix Fig. 11. Consequently, we introduce two more profound challenges in constraints-following assessment.
@@
Q1: How to construct high-quality evaluation data? Many studies focus on evaluating single constraint Chen et al. (2022); Tang et al. (2023), lacking comprehensive analysis across diverse constraints. He et al. (2024b) examines LLM performance on complex real-world instructions but neglect constraint diversity and scenario coverage. Jiang et al. (2023) incrementally incorporate fine-grained constraints to craft multi-level instructions. However, with only 75 instances of mixed type, which risks variability due to limited data, and equating difficulty with constraint quantity oversimplifies the task. Recent work focuses on evaluating constraints combinability Wen et al. (2024). To ensure data quality, we systematically categorize constraints by mining real-world online data and using classification, synthesis, and expert design, covering 10 primary categories and over 25 subcategories. We also cross-match these constraints with various domains and scenarios, ensuring balanced representation and expert-validated reasonableness.
@@
Q2: How to evaluate accurately and meticulously? Evaluating LLMs’ adherence to constraints is challenging and typically involves manual, automated, and programmatic assessments using various metrics. Representative work computes outcomes for verifiable instructions using code  Zhou et al. (2023a); He et al. (2024b).   Jiang et al. (2023) uses scripts and constraint-evolution paths to handle diverse challenging instructions, introducing three metrics tailored to the data’s characteristics. The DRFR method decomposes complex constraints into binary judgments, with GPT evaluating each criterion Qin et al. (2024). Indeed, previous work has ensured the feasibility and objectivity of evaluations through various methods, but they have overlooked assessments from the user’s multiple perspectives. We deconstruct complex instructions from the user’s perspective into multiple sub-needs, categorizing them by priority and constraint type, with LLMs evaluating each checkpoint. Furthermore, a multi-dimensional evaluation criteria is proposed using three metrics from the perspectives of constraints, instructions, and requirements priority.
@@
We introduce CFBench, a comprehensive Chinese benchmark designed to thoroughly evaluate the constraint comprehension and following capabilities of LLMs. CFBench comprises 10 primary categories and over 25 secondary subcategories organized through taxonomic and statistical methodologies. CFBench features 1,000 meticulously curated samples spanning more than 200 real-life scenarios across 20 domains and over 50 NLP tasks, enhancing the breadth and generality of the evaluation data. Additionally, we have seamlessly integrated original instructions and constraint types within each sample, paying particular attention to nuanced combinations, ensuring each constraint is credibly and coherently embedded. Finally, extensive experiments and exploratory discussions provide strong support for evaluation and optimization.
@@
  • We introduce CFBench, a meticulously annotated, large-scale, high-quality Chinese benchmark that encompasses a broad spectrum of real-world scenarios and NLP tasks.
@@
Figure 2: The construction pipeline and evaluation sample of CFBench. Initially, it entails the construction of the constraint system, followed by the assembly of the dataset, and culminating in the proposal of a multi-perspective user view evaluation.
2 Related Work
2.1 Instruction Following
Fine-tuning LLMs with annotated instructional data enhances their ability to follow general language instructions  Weller et al. (2020); Sanh et al. (2022). Studies show that more complex or constrained instructions further improve this ability. For instance, six methods to create intricate instructions from a small set of handwritten seed data are proposed Xu et al. (2023), while  Mukherjee et al. (2023) elevate training data complexity by having GPT-4Achiam et al. (2023) generate reasoning steps for simple instructions. The latest work  Sun et al. (2024a); He et al. (2024a); Dong et al. (2024) suggests that increasing the number and variety of constraints can enhance the complexity of instructions, thereby further improving the model’s ability to follow constraint-based instructions.
2.2 Evaluation of Constraints Following
Constraints such as word count, position, topics, and content have garnered significant attention in the field of Controlled Text Generation  Yao et al. (2023); Zhou et al. (2023b)Zhou et al. (2023a) centers on assessing 25 verifiable instructions. Numerous studies have explored the adherence of LLMs to format constraints, such as complex tabular data  Tang et al. (2023) and customized scenario formats  Xia et al. (2024).  Qin et al. (2024) decomposing a single instruction into multiple constraints. He et al. (2024b) gathered constraints from real-world scenarios and developed a sophisticated benchmark using detailed task descriptions and inputs. Jiang et al. (2023) progressively integrates fine-grained constraints to develop multi-level instructions, thereby enhancing complexity across six distinct types. Concurrent work  Wen et al. (2024), constructs a novel benchmark by synthesizing and refining data from the aforementioned benchmarks, with an emphasis on the combinatorial types of constraints. However, previous studies suffered from fragmented constraints, limited scenarios, and misaligned evaluation methods with user perspectives.
@@ 3 CFBench
3.1 Constraints System
3.1.1 Constraints Collection
We amass a diverse corpus of instructions from real-world scenarios and various NLP tasks  Xia et al. (2024); Li et al. (2024) to ensure a comprehensive system. Initially, we aggregate several million instructions from online logs and NLP tasks, refining these through length filtering and clustering to distill 30,000 high-quality instructions. Utilizing advanced LLM techniques, we extract and expand atomic constraints through evolutionary methods. Using LLMs, we carefully select meaningful atomic constraints, resulting in over 5000 unique constraints.   Domain experts first filter out unreasonable or meaningless atomic constraints and then synthesize these into a structured framework with 10 primary categories and 25 subcategories, guided by principles of statistics, taxonomy, and linguistics.
3.1.2 Constraints System
Content constraints control the scope and depth of output content by specifying certain conditions Zhang et al. (2023), and can be divided into lexical constraints, element constraints, and semantic constraints based on their granularity. Numerical constraints, which ensure that output content meets length and quantity requirements Yao et al. (2023), can be classified into word-level, sentence-level, paragraph-level, document-level based on the objects involved in the planning. Stylistic constraints impart a unique flavor and color to the output, revealing the author’s traits and chosen social objectives Tsai et al. (2021), can be subdivided into tonal, formal, audience, and authorial style constraints based on the perspective of application. Format constraints Tang et al. (2023) standardize expression to guide the generation of complex content and can be categorized into fundamental, bespoke, and specialized scenario constraints based on their usage scenarios.   Linguistic constraints Zhou et al. (2023b) adapt to various scenarios by controlling internal features and logic, grouped into Pragmatic, Syntactic, Morphological, Phonological, and other constraints. Situation constraints Liu et al. (2023) guide response appropriateness through background or situational parameters, can be classified into role-based, task-specific, and complex contextual constraints. Example constraints regulate new responses by leveraging the intrinsic patterns established by a limited set of samples, with an emphasis on assessing the model’s proficiency in contextual constraint learning. Inverse constraints narrow the response space through the mechanism of indirect exclusion.   Contradictory constraints denote conditions that are mutually exclusive, rendering it impossible for the response to fulfill all requirements concurrently, which are prevalent in online logs and are often easily overlooked. Rule constraints define logic flows or actions and meticulously crafted to standardize the road of responses. Details are in Appendix Tab. 8.
3.2 Dataset Construction
@@
3.2.1 Data Source and Selection
Real-world scenarios and NLP tasks form the foundation for CFBench’s initial instructions. By harnessing advanced LLMs, we rigorously assess each instruction for constraint types and quantities within a predefined system, filtering out those with unreasonable or ineffective constraints. Subsequently, we balance the scenarios and constraint types, resulting in a refined set of 2,000 instructions covering all scenarios and NLP tasks. Prompts and checklist generation are in Appendix.
3.2.2 Iterative Refinement
Professional annotators carefully review and refine the data, ensuring the rationality of constraints and gold answer. If modifications are needed, instructions are revised, and LLMs generate responses with refined evaluation criteria, repeating this process until satisfactory results are achieved. Ultimately, comprehensive support is formulated for each sample, detailing high-quality instructions, the ideal answer, specific assessment criteria, constraint types, and priority levels.
@@ 3.2.2 Iterative Refinement
3.3 Dataset Statistics
3.3.1 Overall Statistics
Table 1 provides a statistical overview of CFBench, highlighting substantial differences between the two sets. The Hard Set has more detailed instructions, a greater variety, and a higher number of constraints, indicating higher complexity compared to the Easy Set. The table also shows the diversity and balanced distribution of primary constraint types within CFBench, outperforming other benchmarks. See the Appendix for division details.
3.3.2 Tasks and Domains Distribution
@@
3.3.3 Comparison with Other Benchmarks
As shown in Tab. 3, we thoroughly compare our benchmark with various relevant ones. In terms of size, our benchmark contains approximately twice the number of samples as others. FollowBench Jiang et al. (2023) increases difficulty by adding the number of constraints, but focuses on the incremental increase of a single constraint type. ComplexBench Wen et al. (2024) places more emphasis on the combination relationships between different constraint types, but only designs four types. IFEval Zhou et al. (2023a) focuses on constraints that can be verified, but lacks generalization. Compared to others, CFBench provides comprehensive scenario coverage, diverse systematic constraints, numerous high-quality samples, and multidimensional evaluation. For details, see Appendix Tab. 10, including Case and Features.
3.4 Evaluation Protocol
3.4.1 Evaluation Criteria
We breaking down instructions into multiple simple, independent checkpoints to ensure evaluation accuracy, inspiration was drawn from DRFR Qin et al. (2024). Unlike DRFR, our method emphasizes defining ideal response characteristics and critical evaluation points. The previous sections detailed the checklist generation process, a key part of our evaluation criteria. Furthermore, we employ GPT-4o, as the evaluation model. By repeatedly feeding it the instruction, test model response, and checklist with a carefully tuned prompt, we ensure that the judged response fully meets the judgement format check. This iterative process aims to maximize confidence in our evaluation. The specific evaluation prompt is in the Appendix.
3.4.2 Evaluation Metrics